Prove that the planes 7x +4y -4z +30 =0, 36x - 51y+12z + 17 = 0 14x +8y- 8z -12 = 0 and 12x -17y + 4z -3 =0 form the four faces of a cuboid


1
Expert's answer
2019-09-30T10:59:13-0400

Let plane with equation "7x +4y -4z +30 =0" be "P_1,"

plane with equation

"14x+8y- 8z -12 = 0 \ or\ dividing\ by\ 2:\ 7x+4y-4z-6=0" be plane "P_2,"

plane with equation "36x - 51y+12z + 17 = 0" be "P_3,"

and plane with equation

"12x -17y + 4z -3 =0\ multiplying\ by\ 3:\ 36x-51y+12z-9=0" be plane "P_4."

Let general equation of a plane be "Ax+By+Cz+D=0"

Since planes "P_1" and "P_2" have equal coefficients A, B and C, "P_1" is parallel to "P_2."

Planes "P_3" and "P_4" have equal coefficients A, B and C too, therefor "P_3" is parallel to "P_4."

Normal vector of planes "P_1" and "P_2" is "n_1=(7,4,-4)"

Normal vector of planes "P_3" and "P_4" is "n_2=(36,-51,12)"

Since scalar product "n_1\cdot n_2=7\cdot36+4\cdot(-51)-4\cdot 12=0" ,plane P1 is orthogonal to plane P3.

Plane P2 is parallel to P1 and plane P4 is parallel to P3, therefor P2 is orthogonal to P4.

So, we have:

P1 is parallel to P2,

P3 is parallel to P4,

P1 is orthogonal to P3 and P4,

and P2 is orthogonal to P3 and P4,

therefor these four planes form four faces of cuboid.